OpenStudy (anonymous):

under the QUADRATIC EQUATION(Roots) find two positive real numbers whose product is 16 and whose difference is 1.

OpenStudy (anonymous):

xy = 16 x-y = 1 x = 16/y 16/y - y = 1 16 - y^2 = y y^2 + y -16 = 0 y = 3.53 so x = 4.53 4.53 and 4.53 to 2 decimal places

OpenStudy (anonymous):

check: 3.53 * 4.53 = 15.9909 - close enough!

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

yES!!!

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw